find the maximum value for


y=sin2x+2sin x for all x belong to [ 0, 2 pie]

Grade:

1 Answers

Chetan Mandayam Nayakar
13 years ago

y=sin^2(x)+2sinx=(sinx+1)^2 -1,thus, maximum value=4-1=3